OpenStudy (anonymous):

It takes Sally 6 hours longer than Joan to paint a room. Working together, they can paint the room in 4 hours. How long would it take each person working individually to paint the room?

OpenStudy (saifoo.khan):

joan = x Sally = x+ 6 \[\frac{1}{x} + \frac{1}{x+6} = \frac14\]Solve.:)

OpenStudy (anonymous):

Thanks, so then I get a common denomintor of 24?

OpenStudy (saifoo.khan):

nope, \[\frac{1(x+6) + 1(x)}{x(x+6)} = \frac14\]

OpenStudy (anonymous):

wow I was really off :( All this Algebra jargon is soooooo confusing. Would you mind walking me through the rest of the problem?

OpenStudy (saifoo.khan):

Sure, \[\frac{x+6 + x}{x^2 + 6x } = \frac14\] \[{4(2x+6)}={x^2 + 6x}\]

OpenStudy (anonymous):

ok, so you clear the fraction by multiplying both sides by 4 right? How did you get the bottom equation to the right of the equal sign? I am sorry I am terrible at math and I am confused ;(

OpenStudy (amistre64):

whatever we do to one side has to balance out on the other side. multiply both sides be 4 (x^2+6x) andsimplify

OpenStudy (anonymous):

ok thanks for your help! the whole problem is still very confusing to me.